# Amorphous Boron Nitride as an Ultrathin Copper Diffusion Barrier for Advanced Interconnects

### v2

@@ Abstract
This study focuses on amorphous boron nitride (-BN) as a novel diffusion barrier for advanced semiconductor technology, particularly addressing the critical challenge of copper diffusion in back-end-of-line (BEOL) interconnects. Owing to its ultralow dielectric constant and robust barrier properties, -BN is examined as an alternative to conventional low-k dielectrics. The investigation primarily employs theoretical modeling, using a Gaussian Approximation Potential, to simulate and understand the atomic-level interactions. This machine learning-based approach allows the performance of realistic simulations of amorphous structure of -BN, enabling the exploration of the impact of different film morphologies on barrier efficacy. Furthermore, we studied the electronic and optical properties of the films using a simple Tight-Binding model. In addition to the theoretical studies, we performed diffusion studies of copper through PECVD -BN on Si. The results from both the theoretical and experimental investigations highlight the potential of -BN as a highly effective diffusion barrier, suitable for integration in nanoelectronics. This research shows that -BN is a promising candidate for BEOL interconnects but also demonstrates the synergy of advanced computational models and experimental methods in material innovation for semiconductor applications.
@@ 1 Introduction
The continuous down-scaling of devices and architectures has been the major goal of the electronic industry for decades in order to increase performance, chip density, and reduce power consumption [1, 2, 3]. Interconnect technology is also experiencing significant challenges as devices are scaled down, with sustained efforts to decrease the resistance of metals and to lower the dielectric constant of intermetal dielectrics (IMDs), for optimized control of signal delays such as RC delay.
@@
An exhaustive experimental investigation of the morphology and properties of -BN would be quite time consuming and expensive. On the other hand, atomistic simulations would allow us to study materials at the atomic level and can provide faster insights into the relation between morphology and key parameters of materials with moderate computational cost. Notwithstanding, implementing an accurate modeling of highly disordered materials such as -BN is quite a challenging task. Due to the lack of long-range order, the simulation box of amorphous materials should be quite large. This becomes quickly prohibitive for ab initio calculations. On the other hand, simulations using classical Molecular Dynamics (MD) with empirical potentials can deal with such large samples. However, due to simplifications and assumptions used to develop these interatomic potentials, classical MD cannot capture the diverse atomic environments of the amorphous material. These challenges can be overcome using machine learning-based approaches. Classical MD simulations with machine learning-based interatomic potential trained over ab initio calculations can deal with very large simulation boxes with ab initio accuracy [20, 21].
@@ 2.1 Molecular Dynamics Simulations
MD simulations use Newton’s law of motion to follow the trajectories of atoms through time. The accuracy and reliability of MD simulations largely depends on the accuracy of interatomic potentials. While there are several empirical potentials for -BN and Cu separately, there is no potential that takes both BN and Cu into account. Moreover, empirical potentials provide insufficient accuracy for such complex systems. Machine learning based interatomic potentials that utilize a large dataset of quantum-mechanically computed reference data for structures, energies and forces are quite useful for disordered systems since they enable us to study large numbers of atoms with an accuracy comparable to DFT and a reasonable computing cost. Therefore, we trained a Gaussian Approximation Potential for BNCu systems that is based on DFT-generated energy and force data of a large database. The dataset contains isolated atoms, molecular compounds, crystalline structures, heterostructures and random structures. Details of training and validation of the GAP model is presented in A.
@@
To generate -BN films with different qualities, we employed a melt-quench protocol. This protocol is an efficient way to generate amorphous structures and glasses in MD and has been used extensively by community [18, 19, 22]. In this protocol, an equal number of B and N atoms were placed in a cube simulation box which is periodic in all directions. However, we placed a 25 -thick vacuum region at both the top and bottom of the film to prevent boundary interactions and mimic thin films. After a minimization and an equilibration run at 5000 K for 20 ps with a timestep of 0.25 fs (same timestep used in all simulation), all films were cooled down with a constant cooling rate of 5 - 100 K/ps under NVT ensemble (constant volume-constant temperature) to systematically study the effect of cooling rates on film morphology, capturing variations in the degree of disorder characteristic of amorphous materials. Later all films are relaxed under NVE (constant volume-constant energy) and equilibrated under NVT ensemble for 20 ps each. Here, the initial density of each cube was fixed at approximately 2.1 , the experimentally reported value for -BN films [2]. The number of atoms in 3-nm, 5-nm and 7-nm thick films are respectively 3000, 13000 and 37000 to maintain the selected density. Upon generating the films, several morphological features are investigated to evaluate the quality of samples. We checked the presence of and atoms, homonuclear bonds, nanovoids, as well as the density of the samples. We generated 5 samples at each thickness with each cooling rate to get statistically meaningful data. A 3-nm thick -BN sample generated using this protocol is shown in Fig.1-a.
@@
Later, we investigated the mechanical properties of -BN thin films. Here, we calculate the elastic constants of -BN thin films by computing the stress fluctuations, and the Born matrix of -BN samples [23, 24, 25]. Simulations to compute stress fluctuations and Born matrices were done for 100 ps and data was collected at every 25 fs (100 steps) to get the full elastic constant matrix of each sample. Mechanical properties (Young’s modulus, bulk modulus and shear modulus) were calculated using the Voigt notation [26, 25] which presents the relationship between elastic constants and key mechanical properties.
@@ 3.1 Film Generation and Morphological Features
Fig. 3 presents many interesting trends between cooling rates and morphological features. Fig.3-a shows that higher cooling rates increase the proportion of both B-B and N-N bonds for all thickness. However, the amount of increase in homonuclear bonds varies across film thicknesses. For any given cooling rate, thinner samples consistently have more homonuclear bonds than thicker ones. For instance, the portions of homonuclear bonds in 3-nm and 5-nm thick films generated at a cooling rate of 5 K/ps are approximately 10, while this rises to 30 at higher cooling rates. In contrast, for rapidly cooled (cooling rate 50 K/ps) the 7-nm thick sample, the proportion of homonuclear bonds remains below 20 . Fig. 3-b shows how the coordination number of -BN is influenced by applied cooling rates during film generation. Similarly to what we saw for homonuclear bonds, the effects of cooling rate on the local coordination numbers are more pronounced for thinner films. While all structures cooled at 5 K/ps have no -hybridized atoms (coordination number of 2), their proportion rapidly increases to 10 in 3-nm thick films at higher cooling rates. Similarly, for 5- and 7-nm thick films, the proportions of -hybridized atoms become and , respectively. Moreover, the number of -hybridized atoms (coordination number of 4) drops rapidly at all thicknesses as cooling rates increase. These changes in hybridization are also reflected in the average coordination number of the films, as shown in Fig. 3-c, where larger cooling rates result in a rapid decrease in average coordination for all thicknesses, most notably in the 3-nm thick films. While - and -hybridized fractions show opposing trends with the cooling rate, the average coordination number reflects the overall change in the local atomic environment. With larger cooling rates, the average coordination number of samples at all thicknesses drops quickly, with the largest change observed for 3 nm thick films. A parallel trend is also observed in density of the samples, as depicted in Fig. 3-d. The decrease in coordination number, as cooling rates increase, shows atoms getting far away from each other which leads lower density. While the density of films is for samples cooled with rate of 5 K/ps, the density drops even lower than 2 for 3-nm and 5-nm thick films at higher cooling rate. Meanwhile, even the 7-nm thick samples generated with the largest cooling rate still has an average density of . The significant reduction in density, combined with the increasing number of -hybridized atoms suggests that fast-cooled films have more porous structures. Fig. 11 presented in B shows some voids observed in fast-cooled (with 50 K/ps and 100 K/ps) 3-nm thick -BN films. Around these defects, we also observe a number of -hybridized atoms, which underlines the relationship between the existence of voids and -hybridized atoms.
@@ 3.2 Mechanical Properties
The elastic constants of each film were calculated as explained in the Methods section. Using Voigt notation, Young’s modulus and other moduli were calculated as in Fig. 4. The resulting mechanical properties exhibits parallel trends to the morphological features. Films generated with higher cooling rates, which as seen above present reduced numbers of sp3- and sp2-hybridized atoms, an increased number of homonuclear bonds and higher porosity, display a lower Young’s modulus compared to samples that were cooled slower. Other moduli values are also decreased significantly with higher cooling rates. High Young’s modulus of slow-cooled samples can attributed to high number of - and -hybridized atoms and absence of pores and voids that we observed in fast-cooled samples. The positive impact of -hybridized atoms on mechanical properties is well-established in the literature [18, 19, 37]. Young’s modulus is particularly critical for diffusion barriers, as a higher value indicates greater stiffness and resistance to deformation. This is essential for ensuring that the barrier maintains its structural integrity during the high-stress conditions typical in semiconductor integration and operation. A higher Young’s modulus in slow-cooled films suggests that these barriers would be better at resisting compressive and tensile stresses, minimizing risks such as cracking or delamination that could compromise the diffusion barrier’s effectiveness. In contrast, the lower Young’s modulus observed in fast-cooled films may render them more susceptible to deformation under stress, which could lead to voids and defects during device fabrication and operation. However, it should be noted that high Young’s modulus in barrier layers could be problematic if there is a mismatch between the thermal expansion coefficients of the barrier and other layers. Since high Young’s modulus leads to more stiff structure, such a mismatch could lead to increased stress and structural deformations.
Elastic constants and Voigt notation also allowed us to compute Poisson’s ratio. Although not explicitly shown in the figure, Poisson’s ratio exhibits a weak dependence on cooling rates. For films cooled at 5 K/ps and 10 K/ps, Poisson’s ratio remains relatively consistent, around 0.28–0.30 across different thicknesses, indicating a balance between lateral expansion and contraction under stress. In contrast, faster-cooled samples exhibit a noticeably lower Poisson’s ratio (0.19–0.22), suggesting a stiffer response to deformation. This change, though subtle, may affect the barrier’s ability to adapt to mechanical strains and its performance during thermal cycling and integration processes.
@@ 3.3 Dielectric Properties of -BN Samples
Fig. 5 shows the z-component of the real and imaginary parts of the dielectric function for 3-nm (a, c) and 5-nm thick (b, d) films.444Electronic properties for 7-nm films were not evaluated due to computational limitations. The results are quite similar at both thicknesses. While slow-cooled samples (CR25 K/ps) at both thicknesses exhibited relatively low dielectric constants, around 2.8 (Fig. 5-a), the dielectric functions of the two fastest-cooled samples displayed a tendency to diverge at low frequencies due to the absence of an electronic gap. Although there seems to be a relationship between cooling rates and dielectric constants, the values are too close to establish a definitive trend given the precision of the model.
@@
To get further insight into these transitions and the dielectric constant behaviors, we examine the DOS and IPR of the films. In Fig. 6, we present, for simplicity, the 3-nm thick films generated with cooling rates of 5 K/ps and 100 K/ps. Both films can be seen to exhibit mid-gap states. However, while the slow-cooled sample (6-a) shows a nominal (including mig-gap states) electronic gap of eV, the fast-cooled sample (6-b) displays an effectively closed gap, allowing for the aforementioned low-energy transitions. In the slow-cooled sample, it appears that most midgap states are unoccupied (red dots in Fig. (6-a)), while the fast-cooled sample exhibits occupied mid-gap states. In both cases, mid-gap states show IPRs0, indicating real-space localization, which typically suppresses momentum matrix elements between them (numerators in equations 1 and 1) and DC conductivity, lowering (but here not eliminating) their impact on the dielectric constant.
At this point, we must recall the limitations of our electronic model and give further qualitative remarks. In relation to midgap states, DFT calculations in [28] suggest that the dielectric constant of bulk -BN is increased by the presence of - and - coordinated atoms (in -dominated samples) and low short range order, which can be related to the presence of homonuclear (B-B and N-N) bonds. Lin et al. [15] also point towards the detrimental effects of Boron clusters. Similarly, Glavin et al. [12] indicated that B-B bonds may promote the formation of mid-gap states or conductive pathways in thin films. These remarks should be put in regard with the findings of section 3.1. We further note that although the sample cooled at 25 K/ps had a significant number of N-N bonds, we did not observe any large N or B clusters. However, we observed some B and N clusters (B-B or N-N bonds with a size of 4-12 atoms). The aforementioned morphological features, which we expect in fast-cooled films; especially in the case of thinner films (see Fig. 3) may contribute to the formation of midgap states and significantly affect the dielectric constant in these films. To a degree, we observe such a trend in the electronic calculations presented here, although the simplicity of the tight-binding model used here may underestimate these effects, making quantitative statements difficult. Finally, we should comment on the effects of density: quickly-cooled films; especially the thinnest ones, present a lower density (Fig. 3 (d)), which favorises a lower dielectric constant (see the presence of the unit cell volume in the denominator of equation 1). However, this is accompanied by an increase in undesirable morphological features ( fraction, homonuclear bonds). Within the caveats above, our results suggest that the net effect is not favorable for ultralow dielectric constant properties.

### v3

@@ Abstract
This study focuses on amorphous boron nitride (-BN) as a novel diffusion barrier for advanced semiconductor technology, particularly addressing the critical challenge of copper diffusion in back-end-of-line (BEOL) interconnects. Owing to its ultralow dielectric constant and robust barrier properties, -BN is examined as an alternative to conventional low-k dielectrics. The investigation primarily employs theoretical modelling, using a Gaussian Approximation Potential, to simulate and understand the atomic-level interactions. This machine learning-based approach allows the performance of realistic simulations of amorphous structure of -BN, enabling the exploration of the impact of different film morphologies on barrier efficacy. Furthermore, we studied the electronic and optical properties of the films using a simple Tight-Binding model. In addition to the theoretical studies, we performed diffusion studies of copper through PECVD -BN on Si. The results from both the theoretical and experimental investigations highlight the potential of -BN as a highly effective diffusion barrier, suitable for integration in nanoelectronics. This research shows that -BN is a promising candidate for BEOL interconnects but also demonstrates the synergy of advanced computational models and experimental methods in material innovation for semiconductor applications.
@@ 1 Introduction
The continuous down-scaling of devices and architectures has been the major goal of the electronic industry for decades in order to increase performance, chip density, and reduce power consumption [1, 2, 3]. Interconnect technology is also experiencing significant challenges as devices are scaled down, with sustained efforts to decrease the resistance of metals and to lower the dielectric constant of intermetal dielectrics (IMDs), for optimised control of signal delays such as RC delay.
@@
An exhaustive experimental investigation of the morphology and properties of -BN would be quite time consuming and expensive. On the other hand, atomistic simulations would allow us to study materials at the atomic level and can provide faster insights into the relation between morphology and key parameters of materials with moderate computational cost. Notwithstanding, implementing an accurate modelling of highly disordered materials such as -BN is quite a challenging task. Due to the lack of long-range order, the simulation box of amorphous materials should be quite large. This becomes quickly prohibitive for ab initio calculations. On the other hand, simulations using classical Molecular Dynamics (MD) with empirical potentials can deal with such large samples. However, due to simplifications and assumptions used to develop these interatomic potentials, classical MD cannot capture the diverse atomic environments of the amorphous material. These challenges can be overcome using machine learning-based approaches. Classical MD simulations with machine learning-based interatomic potential trained over ab initio calculations can deal with very large simulation boxes with ab initio accuracy [20, 21].
@@ 2.1 Molecular Dynamics Simulations
MD simulations use Newton’s law of motion to follow the trajectories of atoms through time. The accuracy and reliability of MD simulations largely depends on the accuracy of interatomic potentials. While there are several empirical potentials for -BN and Cu separately, there is no potential that takes both BN and Cu into account. Moreover, empirical potentials provide insufficient accuracy for such complex systems [20, 21, 22]. Machine learning based interatomic potentials that utilize a large dataset of quantum-mechanically computed reference data for structures, energies and forces are quite useful for disordered systems since they enable us to study large numbers of atoms with an accuracy comparable to DFT and a reasonable computing cost. Therefore, we trained a Gaussian Approximation Potential for BNCu systems that is based on DFT-generated energy and force data of a large database. The dataset contains isolated atoms, molecular compounds, crystalline structures, heterostructures and random structures. Details of training and validation of the GAP model are presented in A.
@@
To generate -BN films with different qualities, we employed a melt-quench protocol. This protocol is an efficient way to generate amorphous structures and glasses in MD and has been used extensively by community [18, 19, 22]. In this protocol, an equal number of B and N atoms were placed in a cube simulation box which is periodic in all directions. However, we placed a 25 -thick vacuum region at both the top and bottom of the film to prevent boundary interactions and mimic thin films. After a minimization and an equilibration run at 5000 K for 20 ps with a timestep of 0.25 fs (same timestep used in all simulation), all films were cooled down with a constant cooling rate of 5 - 100 K/ps under NVT ensemble (constant volume-constant temperature) to systematically study the effect of cooling rates on film morphology, capturing variations in the degree of disorder characteristic of amorphous materials. Later all films are relaxed under NVE (constant volume-constant energy) and equilibrated under NVT ensemble for 20 ps each. Here, the initial density of each cube (before the addition of the vacuum regions) was fixed at approximately 2.1 , the experimentally reported value for -BN films [2]. The number of atoms in 3-nm, 5-nm and 7-nm thick films are respectively 3000, 13000 and 37000 to maintain the selected density. Upon generating the films, several morphological features are investigated to evaluate the quality of samples. We checked the presence of and atoms, homonuclear bonds, nanovoids, as well as the density of the samples. We generated 5 samples at each thickness with each cooling rate to get statistically meaningful data. A 3-nm thick -BN sample generated using this protocol is shown in Fig.1-a.
@@
Later, we investigated the mechanical properties of -BN thin films. The elastic constants of -BN thin films were calculated by computing the stress fluctuations and the Born matrix of -BN samples [23, 24, 25]. These simulations were performed over 100 ps, with data collected every 25 fs (100 steps), to obtain the complete elastic constant matrix for each sample. Mechanical properties (Young’s modulus, bulk modulus, and shear modulus) were then derived from the elastic constants obtained in the MD simulations [26, 25].
@@ 3.1 Film Generation and Morphological Features
Fig. 3 presents many interesting trends between cooling rates and morphological features. Fig.3-a shows that higher cooling rates increase the proportion of both B-B and N-N bonds for all thickness. However, the amount of increase in homonuclear bonds varies across film thicknesses. For any given cooling rate, thinner samples consistently have more homonuclear bonds than thicker ones. For instance, the portions of homonuclear bonds in 3-nm and 5-nm thick films generated at a cooling rate of 5 K/ps are approximately 10, while this rises to 30 at higher cooling rates. In contrast, for rapidly cooled (cooling rate 50 K/ps) the 7-nm thick sample, the proportion of homonuclear bonds remains below 20 . Fig. 3-b shows how the coordination number of -BN is influenced by applied cooling rates during film generation. Similarly to what we saw for homonuclear bonds, the effects of cooling rate on the local coordination numbers are more pronounced for thinner films. While all structures cooled at 5 K/ps have no -hybridized atoms (coordination number of 2), their proportion rapidly increases to 10 in 3-nm thick films at higher cooling rates. Similarly, for 5- and 7-nm thick films, the proportions of -hybridized atoms become and , respectively. Moreover, the number of -hybridized atoms (coordination number of 4) drops rapidly at all thicknesses as cooling rates increase. These changes in hybridization are also reflected in the average coordination number of the films, as shown in Fig. 3-c, where larger cooling rates result in a rapid decrease in average coordination for all thicknesses, most notably in the 3-nm thick films. While - and -hybridized fractions show opposing trends with the cooling rate, the average coordination number reflects the overall change in the local atomic environment. With larger cooling rates, the average coordination number of samples at all thicknesses drops quickly, with the largest change observed for 3 nm thick films. A parallel trend is also observed in density of the samples,444Like for the dielectric function, we used the difference in z-coordinate between the atoms of highest and lowest and the area of the periodic cell to estimate the film volume . as depicted in Fig. 3-d. The decrease in coordination number, as cooling rates increase, shows atoms getting far away from each other which leads lower density. While the density of films is for samples cooled with rate of 5 K/ps, the density drops even lower than 2 for 3-nm and 5-nm thick films at higher cooling rate. Meanwhile, even the 7-nm thick samples generated with the largest cooling rate still has an average density of . The significant reduction in density, combined with the increasing number of -hybridized atoms suggests that fast-cooled films have more porous structures. Fig. 11 presented in B shows some voids observed in fast-cooled (with 50 K/ps and 100 K/ps) 3-nm thick -BN films. Around these defects, we also observe a number of -hybridized atoms, which underlines the relationship between the existence of voids and -hybridized atoms.
The observed reduction in homonuclear bonds and -hybridized atoms with increasing thickness, as shown in Fig. 3, suggests enhanced structural relaxation in thicker films. Interestingly, this improvement in structural order becomes more pronounced in the 7-nm films compared to the 3-nm and 5-nm films, where the differences are relatively smaller. This significant improvement in the 7-nm films reflects the ability to better accommodate strain and achieve a more ordered atomic arrangement as thickness increases. Similar behaviour has been reported in the literature for amorphous systems [35, 36, 37], where larger volumes facilitate relaxation processes and reduce disorder.
@@ 3.2 Mechanical Properties
The elastic constants of each film were calculated as explained in the Methods section. Young’s modulus and other mechanical moduli of films respect to the thickness and applied cooling rate is presented in Fig. 4. The resulting mechanical properties exhibit trends that parallel the morphological features. Films generated with higher cooling rates, which as seen above present reduced numbers of - and -hybridized atoms, an increased number of homonuclear bonds and higher porosity, display a lower Young’s modulus compared to samples that were cooled slower. Other moduli values are also decreased significantly with higher cooling rates. High Young’s modulus of slow-cooled samples can attributed to high number of - and -hybridized atoms and absence of pores and voids that we observed in fast-cooled samples. The positive impact of -hybridized atoms on mechanical properties is well-established in the literature [18, 19, 40]. Young’s modulus is particularly critical for diffusion barriers, as a higher value indicates greater stiffness and resistance to deformation. This is essential for ensuring that the barrier maintains its structural integrity during the high-stress conditions typical in semiconductor integration and operation. A higher Young’s modulus in slow-cooled films suggests that these barriers would be better at resisting compressive and tensile stresses, minimizing risks such as cracking or delamination that could compromise the diffusion barrier’s effectiveness. In contrast, the lower Young’s modulus observed in fast-cooled films may render them more susceptible to deformation under stress, which could lead to voids and defects during device fabrication and operation. However, it should be noted that high Young’s modulus in barrier layers could be problematic if there is a mismatch between the thermal expansion coefficients of the barrier and other layers. Since high Young’s modulus leads to more stiff structure, such a mismatch could lead to increased stress and structural deformations.
Although not explicitly shown in the figure, Poisson’s ratio exhibits a weak dependence on applied cooling rates. For films cooled at 5 K/ps and 10 K/ps, Poisson’s ratio remains nearly constant at approximately 0.28–0.30 across different thicknesses, suggesting a consistent balance between lateral expansion and contraction under stress. In contrast, faster-cooled samples exhibit a noticeably lower Poisson’s ratio (0.19–0.22), suggesting a stiffer response to deformation. This change, though subtle, may affect the barrier’s ability to adapt to mechanical strains and its performance during thermal cycling and integration processes.
@@ 3.3 Dielectric Properties of -BN Samples
Fig. 5 shows the z-component of the real and imaginary parts of the dielectric function for 3-nm (a, c) and 5-nm thick (b, d) films.555Electronic properties for 7-nm films were not evaluated due to computational limitations. The results are quite similar at both thicknesses. While slow-cooled samples (CR25 K/ps) at both thicknesses exhibited relatively low dielectric constants, around 2.8 (Fig. 5-a), the dielectric functions of the two fastest-cooled samples displayed a tendency to diverge at low frequencies due to the absence of an electronic gap. Although there seems to be a relationship between cooling rates and dielectric constants, the values are too close to establish a definitive trend given the precision of the model.
@@
To get further insight into these transitions and the dielectric constant behaviours, we examine the DOS and IPR of the films. In Fig. 6, we present, for simplicity, the 3-nm thick films generated with cooling rates of 5 K/ps and 100 K/ps. Both films can be seen to exhibit mid-gap states. However, while the slow-cooled sample (6-a) shows a nominal (including mig-gap states) electronic gap of eV, the fast-cooled sample (6-b) displays an effectively closed gap, allowing for the aforementioned low-energy transitions. In the slow-cooled sample, it appears that most midgap states are unoccupied (red dots in Fig. (6-a)), while the fast-cooled sample exhibits occupied mid-gap states. In both cases, mid-gap states show IPRs0, indicating real-space localization, which typically suppresses momentum matrix elements between them (numerators in equations 1 and 2) and DC conductivity, lowering (but here not eliminating) their impact on the dielectric constant.
At this point, we must recall the limitations of our electronic model and give further qualitative remarks. In relation to midgap states, DFT calculations in [28] suggest that the dielectric constant of bulk -BN is increased by the presence of - and - coordinated atoms (in -dominated samples) and low short range order, which can be related to the presence of homonuclear (B-B and N-N) bonds. Lin et al. [15] also point towards the detrimental effects of Boron clusters. Similarly, Glavin et al. [12] indicated that B-B bonds may promote the formation of mid-gap states or conductive pathways in thin films. These remarks should be put in regard with the findings of section 3.1. We further note that although the sample cooled at 25 K/ps had a significant number of N-N bonds, we did not observe any large N or B clusters. However, we observed some B and N clusters (B-B or N-N bonds with a size of 4-12 atoms). The aforementioned morphological features, which we expect in fast-cooled films; especially in the case of thinner films (see Fig. 3) may contribute to the formation of midgap states and significantly affect the dielectric constant in these films. To a degree, we observe such a trend in the electronic calculations presented here, although the simplicity of the tight-binding model used here may underestimate these effects, making quantitative statements difficult. Finally, we should comment on the effects of density: quickly-cooled films; especially the thinnest ones, present a lower density (Fig. 3 (d)), which favours a lower dielectric constant (see the presence of the unit cell volume in the denominator of equation 1). However, this is accompanied by an increase in undesirable morphological features ( fraction, homonuclear bonds). Within the caveats above, our results suggest that the net effect is not favourable for ultralow dielectric constant properties.